OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6.

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

y + 3 = -6(x + 2) y + 2 = 6(x + 3) y - 3 = -6(x - 2) y - 2 = 6(x - 3) those are the options

OpenStudy (anonymous):

PLEASE

OpenStudy (anonymous):

id go with the 3rd one

OpenStudy (anonymous):

you sure?

OpenStudy (anonymous):

can you help me out with another question? :) pwease

OpenStudy (anonymous):

?

OpenStudy (anonymous):

ill see what I can do

OpenStudy (anonymous):

omg thank you.....

OpenStudy (anonymous):

Identify the correct x intercept and y intercept of the equation x + 7y = -21 A) x intercept = (3, 0); y intercept = (0, 21) B) x intercept = (-3, 0); y intercept = (0, -21) C) x intercept = (21, 0); y intercept = (0, 3) D) x intercept = (-21, 0); y intercept = (0, -3)

OpenStudy (anonymous):

not suree on this one

OpenStudy (anonymous):

Choose the equation of the line passing through the point (6, -2) and parallel to y = x + 7. A) y = x + 8 B) y = x - 4 C) y = x + 4 D) y = x - 8 what about this one?

OpenStudy (anonymous):

?????????????

OpenStudy (anonymous):

said I wasn't suree

OpenStudy (anonymous):

that was a diff one

OpenStudy (anonymous):

nevermind. I gotta go